An object is launched straight up into the air at an initial velocity of 64 feet per second. It is launched from an initial height of 6 feet off the ground. Find all times t that the object is at a height of 54 feet off the ground.



Answer by ankor@dixie-net.com(22740)

An object is launched straight up into the air at an initial velocity of 64 feet per second.
It is launched from an initial height of 6 feet off the ground.
Find all times t that the object is at a height of 54 feet off the ground.
:
Using the formula h = -16t^2 + vt + c, where
h = the height above ground at t time
t = time since initially launched
v = initial velocity
c = height above ground at launch
:
h = 54
-16t^2 + 64t + 6 = 54
-16t^2 + 64t + 6 - 54 = 0
-16t^2 + 64t - 48 = 0
simplify, divide by -16
t^2 - 4t + 3 = 0
factors to
(t-3)(t-1) = 0
two solutions
t = 1 sec at 54 ft on the way up
and
t = 3 sec at 54 ft on the way down